1/m+1/n = 1/t solve for t
Answer by ewatrrr(24785)


Hi,
1/m+1/n = 1/t solve for t
Mulitply each term on both sides of the equation by mnt to clear the denominators
nt + mt = mn
t(n+m) = mn
t = mn/(n+m)
.
If we were to be solving for m, we would isolate m on one side of the equation
nt + mt = mn
mt- mn = -nt
m(t -n) = -nt
m = -nt/(t-n)